OpenStudy (anonymous):

Determine is the following is an inner product space on R^2

OpenStudy (anonymous):

\[A=\left(\begin{matrix}a_{1} \\ a_{2}\end{matrix}\right)\] \[B=\left(\begin{matrix}b_{1} \\ b_2\end{matrix}\right)\] A dot B= \[a_1b_1-a_2b_1-a_1b_2+3_2b_2\]

OpenStudy (fibonaccichick666):

yes? and

OpenStudy (anonymous):

I'm having trouble verifying clause 1= A dot B=B dot A

OpenStudy (fibonaccichick666):

ohh ok h/o

OpenStudy (anonymous):

I...don't know. The equation I wrote is the defined dot product of A dot B

OpenStudy (fibonaccichick666):

ok let's see if we can reverse engineer it

OpenStudy (fibonaccichick666):

what is the det(AB) (the matrix composed of the two?

OpenStudy (anonymous):

Can there be a determinant of a non-square matrix?

OpenStudy (fibonaccichick666):

\[ \left [AB \right] = \left [a_1,b_1;a_2,b_2 \right]^T\]

OpenStudy (anonymous):

???

OpenStudy (fibonaccichick666):

that's the matrix I was talking about. The one composed of the two

OpenStudy (fibonaccichick666):

\[\left[\begin{matrix}a_1 & b_1 \\ a_2 & b_2\end{matrix}\right]\]

OpenStudy (anonymous):

Ahh. Determinant would be... a1b2-a2b1 yes?

OpenStudy (fibonaccichick666):

yea kind of reminds you of their dot product right

OpenStudy (fibonaccichick666):

also that 3 is that an a up there?

OpenStudy (anonymous):

yes, but it's missing some terms from the defined dot product. Also yes, last term is 3a2b2

OpenStudy (fibonaccichick666):

ok the 3... I'm not sure how to reverse that this method won't help give me a min to think about the new problem

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

I should stop you here and say that I've now verified clauses 1,2, and 3. I'm having trouble proving clause 4 to be greater than or equal to 0

OpenStudy (fibonaccichick666):

so lets see we can write \(A \dot\ B\) as: \[1(a_1b_1)-1(a_2b_1)-1(a_1b_2)+3(a_2b_2)\] right?

OpenStudy (anonymous):

Sorry, been working.

OpenStudy (fibonaccichick666):

lol it's ok

OpenStudy (fibonaccichick666):

so try doing \(A \dot\ A\) see what you get

OpenStudy (anonymous):

Clause 4 states that V dot V is greater than or equal to zero. How do I prove that V dot V is greater than or equal to zero for all cases? I'm about to check if it's for all real V, or V greater than zero or whatnot.

OpenStudy (anonymous):

A dot A gives me \[a_1^{2}-2a_2a_1+3a_2^{2}\]

OpenStudy (fibonaccichick666):

and can you factor that?

OpenStudy (anonymous):

Umm.

OpenStudy (anonymous):

(a1+a2)(a1-3a2)?

OpenStudy (fibonaccichick666):

also clause 4 is that d(A,B)≥0 for me is that yours too? with it =0 iff A=B

OpenStudy (anonymous):

I have a dot A is greater than or equal to zero. and A dot A is equal to zero if, and only if v=0

OpenStudy (anonymous):

Also, my above doesn't factor. The last term becomes negative in my factoring

OpenStudy (fibonaccichick666):

hmm interesting. on the property and yea if it doesn't factor what could that imply?

OpenStudy (anonymous):

That...it doesn't equal zero?

OpenStudy (fibonaccichick666):

ok so then by mine it is not an IPS but hmm let me see if I can factor it and look at your def.

OpenStudy (fibonaccichick666):

I'm going to bring in some backup just in case. @e.mccormick can you take a look at this?

OpenStudy (anonymous):

Hey mccormick, thanks for coming.

OpenStudy (e.mccormick):

/wave Looks like you are having fun.

OpenStudy (fibonaccichick666):

Of course, Linear Algebra is just AAMAzing *rolls eyes*

OpenStudy (anonymous):

Of course. Now, I can state that clause four is true for specific cases, but how do I say that's it's true for all v?

OpenStudy (e.mccormick):

Well, for it to be an inner product, the result must be nonnegative. That is one thing.

OpenStudy (anonymous):

Well yes. But how do I express that? The equation I got for A dot A is above

OpenStudy (fibonaccichick666):

yes because as per my Clause 4 \(\sqrt{A\dot\ A}=0\)

OpenStudy (anonymous):

My clause four states: A dot A is greater than or equal to zero, AND A dot A is equal to zero if, and only if, v=0

OpenStudy (fibonaccichick666):

hmm... could we use both?

OpenStudy (fibonaccichick666):

oh wait your 4 is my 1 sorry found it

OpenStudy (fibonaccichick666):

no wonder i was confused

OpenStudy (e.mccormick):

So, you are basically proving the four axioms of inner products, which are essentially the same as dot products, for this particualr case, right?

OpenStudy (fibonaccichick666):

yep but with an odd definition of dot product. It's a weighted Euclidean inner product like problem

OpenStudy (anonymous):

Right. With that weird definition on the dot product, I can't say that A dot A= A^2 and therefore greater than or equal to zero

OpenStudy (e.mccormick):

basically you need to confirm all four, or break one. If one is a problem to test, move on and save it for last. See if any of the other three break.

OpenStudy (fibonaccichick666):

b=1/3 (a-I Sqrt[2] a) but because it has an imaginary root that means it is negative

OpenStudy (fibonaccichick666):

right?

OpenStudy (fibonaccichick666):

he already did the other 3

OpenStudy (e.mccormick):

Ah.... well... I did skip to the end here. LOL

OpenStudy (fibonaccichick666):

lol cheater ;P

OpenStudy (anonymous):

I don't understand your answer Fibonacci. There is no B in this problem, just A and A.

OpenStudy (fibonaccichick666):

oops change the b to \(a_2\)

OpenStudy (anonymous):

Right, but does it matter if a2 is negative? It's the dot product that cannot be negative, not one specific term.

OpenStudy (anonymous):

Though it gives me a thought

OpenStudy (anonymous):

Scratch my thought. Anyone else got one?

OpenStudy (fibonaccichick666):

\[a_2=\frac{1}{3}(a-ia\sqrt{2})~and~a_2=\frac{1}{3}(a-ia\sqrt{2})\]

OpenStudy (anonymous):

I don't quite understand the reasoning. Wait, if both a1 and a2 do not exist, then a1^2 is negative, and so is a2^2. And a1a2 is also negative. Therefore...

OpenStudy (fibonaccichick666):

nah there are 2 \(a_2\)

OpenStudy (anonymous):

So how would I express that?

OpenStudy (fibonaccichick666):

well I think I would go for the \(A \dot\ A =0\) clause if you have an imaginary number that isn't in \( R^2\)

OpenStudy (anonymous):

???

OpenStudy (anonymous):

Both parts of clause 4 have to be true

OpenStudy (fibonaccichick666):

If \(A \dot\ A=0\) at the points I mentioned above then 1) A is not contained in \(R^2\) and 2) \(A \not=0\)

OpenStudy (anonymous):

I still don't understand. How would you write out clause 4 from start to finish?

OpenStudy (fibonaccichick666):

Well to start, No, it is not an inner product space for \(R^2\) because...

OpenStudy (fibonaccichick666):

do you agree?

OpenStudy (e.mccormick):

So FibonacciChick666, you are saying it is not an inner product because non-zero vectors can have a norm of zero?

OpenStudy (fibonaccichick666):

yes?... I'm confusing myself now

OpenStudy (e.mccormick):

I think so, because that would invatidate it.

OpenStudy (fibonaccichick666):

it's by contradiction surely

OpenStudy (anonymous):

I'm lost

OpenStudy (fibonaccichick666):

ok so have you had a proof writing class?

OpenStudy (anonymous):

No, unfortunately. I did proofs back in geometry about 4 years ago, and we've done some proofs in linear algebra. No formal proof course though

OpenStudy (fibonaccichick666):

okk that makes this especially hard(i just had the same problem this semester) so what I'm sure e.mccormick is typing is a proof

OpenStudy (fibonaccichick666):

we will guide you through said proof ok?

OpenStudy (e.mccormick):

If there is a way to make \(a_2>a_1\) and negative when it comes out of the root, which I think the \(a_2=\frac{1}{3}(a-ia\sqrt{2})~and~a_2=\frac{1}{3}(a-ia\sqrt{2})\) is showing, then you get 3 times a negative, and larger than the rest of the \(a_1^{2}-2a_2a_1+3a_2^{2}\). Wouldn't that break a few things?

OpenStudy (anonymous):

Alright. After this question I've got one more for the board tonight. I can type it here for you two if you'd like to help (it's on cross product proofs) or I can make a new question.

OpenStudy (fibonaccichick666):

new question would be preferable due to the long list we currently have going :)

OpenStudy (anonymous):

ok, but after we figure this out. Now, Mccormick. I have to prove a2 larger than a1 (or smaller, if negative) for this to fail?

OpenStudy (fibonaccichick666):

If we just assume A>0 for the start. WTS \(A \dot\ A\not=0\)

OpenStudy (fibonaccichick666):

plug and chug

OpenStudy (anonymous):

If I could just assume for a proof like this i wouldn't need to work on the problem

OpenStudy (fibonaccichick666):

then get some values where it = 0

OpenStudy (anonymous):

How?

OpenStudy (fibonaccichick666):

we can assume because we will show the other in another step

OpenStudy (anonymous):

That doesn't make sense to me

OpenStudy (anonymous):

This would be easiest if I could come up with a number for a1 and a number for a2 such that (a1^2)+2(a1)(a2)+)a2^2)=0 where a1=/=0 or a2=/=0

OpenStudy (fibonaccichick666):

well actually we wouldn't need another step because it violated our initial condition

OpenStudy (e.mccormick):

Inductive proofs make an assumption and either confirm it or prove it wrong. If they confirm it, it is a proof. it they deny it it is a proof by contradiction that eliminates the posibility.

OpenStudy (anonymous):

meaning?

OpenStudy (e.mccormick):

You are allowed to assume things. You just need to catch up to what was assumed or blow it away.

OpenStudy (anonymous):

So if I let A be greater than or equal to zero, how would I express that as a step and have it make sense?

OpenStudy (fibonaccichick666):

Assume A>0